What is the pH of a 15.9 M solution of H2SO4? The first proton completely dissociates; the Ka for the second proton is 1.2×10–2.
Answer #1

Following is the - complete Answer -&- Explanation: for the given: Question: in...typed format...

\RightarrowAnswer:

For the given solution of H2SO4 (aq) : pH = 1.92 , neglecting any negative ( - )ve value of : pH ... ( second dissociation)

\RightarrowExplanation:

Following is the complete: Explanation: for the above: Answer...

  • Given:
  1. Molar concentration of H2SO4 (aq) : Macid = 15.9 M  ( mol/L ) ...i.e. assumed = Z (mol/L )
  2. Acid dissociation constant for the second proton: Ka2 = 1.2 x 10-2
  3. For the solution of H2SO4 (aq) : first proton: completely dissociates.
  • ​​​​​​​Step - 1:

​​​​​​​For dissociation of aqueous solution of H2SO4 (aq) , we can form the following sequence of balanced chemical equations:

Type of dissociation Balanced chemical equation Identity of Equation
First proton dissociation H2SO4 (aq)  \rightleftharpoons H+(aq) + HSO4- (aq) Equation - 1
Second proton dissociation HSO4- (aq) \rightleftharpoons H+(aq) + SO42- (aq) Equation - 2
  • Step - 2:

​​​​​​​\Rightarrow Since, the dissociation of the first proton: was completed, according to: Equation -1, we can assume the concentration: of the proton , i.e. [ H+ ], after the first dissociation: remains constant: at:  [ H+ ] = Z ( mol/L )

Therefore, after dissociation of the second proton: let's assume, the concentration of : [ H+ ] , will become the following:

i.e. after dissociation of the second proton: \Rightarrow[ H+ ] = Z + x   ( mol/L )

\Rightarrow Therefore: we will have the following molar concentrations: ( i.e. after the dissociation of second proton ) ;

  1. [ H+ ] = Z + x   ( mol/L )
  2. [HSO4- ] = Z - x ( mol/L )
  3. [SO42- ] = x   ( mol/L )

​​​​​​​Note: We have arrived at the above results, according to the following ICE Table, prepared based on: Equation - 2...

  • Step - 3:

​​​​​​​Let's assume, initial concentration: of HSO4- (aq) = Z ( mol/L ) = 15.9 M ( given )

Therefore: following will be our ICE Table:

[HSO4- ] ,M (mol/L) [ H+ ], M (mol/L )   [SO42- ], M (mol/L )
Initial concentration Z 0.0 0.0
Change ( in concentration ) Z - x +x +x
Equilibrium ( concentration ) Z - x +x +x

\RightarrowTherefore:

we will have the following molar concentrations: ( i.e. after the dissociation of second proton, at Equilibrium ) ;

  1. [ H+ ]eq = Z + x   ( mol/L )
  2. [HSO4- ]eq = Z - x ( mol/L )
  3. [SO42- ]eq  = x   ( mol/L )
  • Step - 4:

Therefore: we will get:

\RightarrowKa2 =   ( [ H+ ]eq x [SO42- ]eq ) / (  [HSO4- ]eq ) =   1.2 x 10-2   

Plugging in values, in the above: Equation: we will get the following:

\Rightarrow   Ka2 = 1.2 x 10-2   = [ ( Z + x ) x ( x ) ] / ( Z - x ) ----------------------Equation - 3

\RightarrowWe know:  Z = 15.9 M ( mol/L )

Plugging in values in, Equation - 3, we will get the following:

  • Step - 5:

​​​​​​​\Rightarrowx2 + 15.912 x - 0.1908 = 0 -----------------Equation - 4

By solving the above: Quadratic Equation; and neglecting: (negative values of 'x' , we will get the following:

\Rightarrowx = 0.01198 M ( mol/L ), approx.

Therefore:

  • Step - 6:

​​​​​​​Therefore: we get...

\Rightarrow [ H+ ]eq = Z + x = 15.9 + 0.01198 = 15.91198 M   ( mol/L )

Therefore:

\Rightarrow   pH = - log10 [ H+ ]eq =   - log10 [ 15.91198 ] =  -1.20

OR, if we assume:   [ H+ ]eq = x = 0.01198 M

Then, we will have the following:

'\Rightarrow    pH = - log10 [ H+ ]eq =   - log10 [ 0.01198 ] =  1.92

\Rightarrow   pH = 1.92